# Subsystem Information Capacity in Random Circuits and Hamiltonian Dynamics

### v2

@@ III.2.6 Random Floquet circuit with many-to-many encoding
In one-dimensional random Clifford Floquet circuits of Anderson localization phases, the mutual information is observed to be localized around the $L/2$ segment $E$ as shown in Fig. 18. Again, consistent with Anderson localization results with one-to-one encoding, in the thermodynamic limit, the information will be fully localized in the $L_{A}=L/2$ subsystem, and the steady state SIC gives $2$ for any $L_{A}/L>1/2$.
@@
Figure 18: Steady-state SIC in 1D random Clifford Floquet circuit with different subsystem sizes $L_{A}$ under finite rate encoding scheme that encodes $L/2$ Bell pairs with $L/2$ ancilla qubits at the beginning. The results are average over 400 configurations.
IV Results for Hamiltonian dynamics
@@
In this section, we shift our focus to Hamiltonian quench dynamics described by one-dimensional fermionic systems. We investigate information dynamics in non-equilibrium phases utilizing AA model and SSH model, respectively. Unlike the random circuit case, where the applied local unitaries are generic without any symmetry or conservation laws, the Hamiltonians investigated in this section are U(1) symmetric and thus follow charge conservation. U(1)-symmetric evolution often gives rise to a spectrum of novel phenomena that are distinct from those in generic evolution without any conservation laws. Examples include OTOC Khemani et al. (2018); Rakovszky et al. (2018), Rényi entanglement growth Rakovszky et al. (2019); Huang (2020); Zhou and Ludwig (2020); Žnidarič (2020), symmetry restoration Liu et al. (2024c), and measurement-induce charge sharpening transitions Agrawal et al. (2022).
@@
In this section, we focus on the one-to-one encoding scheme and von Neumann entropy based mutual information. Notably, fermionic GHZ states generated by one-to-all encoding are beyond Gaussian states and thus cannot be simulated by Gaussian state simulators for large-size systems even the evolution Hamiltonian itself is quadratic (see Appendix D). It is worth noting that in U(1) symmetric systems, Rényi entropy can show different dynamical behaviors (diffusive growth) than the von Neumann entropy (ballistic growth). We leave the information dynamics described by Rényi mutual information and Holevo information as an interesting future direction to explore.
IV.1 Quasiparticle picture for SIC dynamics in non-interacting Hamiltonians
@@
The entanglement growth in noninteracting and integrable systems following a quench can be successfully explained by the intuitive quasiparticle picture Calabrese and Cardy (2005); Alba and Calabrese (2017); Calabrese (2020); Orito and Imura (2023). The quasiparticles of the quenched Hamiltonian are formed in pairs or multiplets and uniformly distributed in space. These quasiparticles go ballistically with the speed determined by the dispersion relations and there is no interaction between different momentum species of quasiparticles. Specifically, only quasiparticle pairs of momentum ±$\pm k$ with one residing on intervals $A$ and the other residing on its complementary $\bar{A}$ contribute to the entanglement entropy $S_{A}$ one unit $s_{k}$. Here $s_{k}$ is the entropy unit $s_{k}=-n_{k}\log n_{k}-(1-n_{k})\log(1-n_{k})$ where $n_{k}$ represents the quasiparticle density for momentum $k$. For Néel initial states and the XX Hamiltonian which is equivalent to free fermion hopping model, $n_{k}=1/2$ for all quasiparticle momentum $k$ with the velocity for quasiparticle as $v_{k}=\partial\varepsilon_{k}/\partial k=-2\sin k$ and entropy unit $s_{k}=1$ in the $\log_{2}$ base. The introduction of onsite disorder renders fermion operators of different momentum coupled, which is evident in Fourier transformed Hamiltonian in momentum space as $\sum_{k,k^{\prime}}\mu(k-k^{\prime})c^{\dagger}_{k}c_{k^{\prime}}$. Therefore, the onsite disorder leads to strong scattering between quasiparticles of different momentum, slowing down and invalidating the ballistic transport of quasiparticles.
@@
IV.2 Results for non-interacting Hamiltonians
In this subsection, we focus on the results of the non-interacting AA model ($U=0$). The model in Eq. (1) has an analytical exact critical point separating the Anderson localization and extended phases at $w_{c}=J$ Harper (1955); Aubry and André (1980).
In the absence of quasiperiodic potential, i.e. $w=0$, the results can be interpreted by the quasiparticle picture as described above. As shown in Fig. 19(b), the theoretical prediction based on the quasiparticle picture is qualitatively in agreement with the numerical results. The deviating time from the $MI=2$ plateau happens at $t^{*}=L_{A}/(2v_{max})$ where $v_{max}$ is the maximal speed of quasiparticles as $\text{max}_{k}\;2\sin k=2$. After $t^{*}$, the fastest quasiparticle from $E$ begins to leave the interval $A$, leading to the decline in SIC. Additionally, the finite size system with periodic boundary conditions renders the oscillation of SIC dynamics in the clean limit $w=0$ when the quasiparticle picture is valid without scattering between different species. The oscillation behavior is similar as reported in Modak et al. (2020) where the quasiparticles can go around the system and get back to the interval $A$ in late times. It is important to note that the presence and absence of oscillation behaviors and the exact periodicity for the oscillation both depend on the quasiparticle dispersion relations $\epsilon_{k}$. Our results here are determined by the free fermion dispersion $\epsilon_{k}\propto\cos k$. Investigating the (non)oscillation behaviors for other dispersion relations presents a promising avenue for future research.
@@ IV.2 Results for non-interacting Hamiltonians
The late-time steady-state behavior for the SIC $I(A:R)$ is given in Fig. 21. For the clean limit $w=0$, the late-time SIC forms a straight diagonal line, indicating a uniform spatial distribution of the initial information at late times. A slightly larger $w$ will make the line bend towards the upper left. After reaching Anderson localization phases $w>w_{c}$, the SIC maintains the value $2$ for any $L_{A}/L>0$ in the thermodynamic limit. The curves for different system sizes converge in Anderson localization phases as shown in the inset of Fig. 21, reflecting the system size-independent nature of the localization length. This is similar to the Anderson localization phase in the random Clifford Floquet circuit.
@@
IV.3 Results for interacting Hamiltonians
@@
IV.3.1 Thermal and many-body localization systems
Many-body localization is a novel non-equilibrium phase that challenges the eigenstate thermalization hypothesis Srednicki (1994). Building upon Anderson localization Anderson (1958), the system can evade thermalization with many-body interactions Basko et al. (2006). A variety of mechanisms are proposed to stabilize MBL phases, including strong random disorder Oganesyan and Huse (2007); Pal and Huse (2010); Altman and Vosk (2015); Nandkishore and Huse (2015); Abanin et al. (2019), quasiperiodic potentials Iyer et al. (2013); Khemani et al. (2017); Zhang and Yao (2018, 2019); Kohlert et al. (2019) and linear potentials Schulz et al. (2019); van Nieuwenburg et al. (2019); Khemani et al. (2020); Doggen et al. (2021); Liu et al. (2023b). As a fundamental basis for other emergent dynamical phases such as discrete time crystal Else et al. (2020); Zaletel et al. (2023) and Hilbert space fragmentation Khemani et al. (2020); Doggen et al. (2021); Yang et al. (2020), MBL has been extensively explored on various experimental platforms of programmable quantum simulators and quantum computers Schreiber et al. (2015); Smith et al. (2016); Liu et al. (2023c); Gong et al. (2021). In this work, we explore MBL phases through the lens of information protection and information dynamics.
Numerous studies have explored information scrambling in MBL systems, mainly using the probe of OTOC Fan et al. (2017); Swingle and Chowdhury (2017); Huang et al. (2017); He and Lu (2017); Chen et al. (2017) and TMI Bölter and Kehrein (2022); MacCormack et al. (2021); Orito et al. (2022). However, these studies have not quantitatively determined the late-time redistribution of information from local perturbations. In this work, we use the newly introduced subsystem information capacity as the probe to investigate the information dynamics in MBL systems. The results are anticipated to be very different from Anderson localization cases since entanglement growth in the two phases is very different (no growth vs. logarithmic increase) Bardarson et al. (2012); Serbyn et al. (2013).
@@ IV.3.1 Thermal and many-body localization systems
The phase diagram of the interacting AA model is determined in Appendix E with $w_{c}\approx 1.4\pm 0.2$ for MBL transitions in our parameter regimes. The dynamical and steady-state SIC for different subsystem sizes $L_{A}$ and different potential strengths $w$ are depicted in Fig. 22 and 23, respectively. Regarding the dynamics, we observe a logarithmic decay in SIC in MBL phases, rendering a very slow saturation for the subsystem SIC as shown in Fig. 22(b). Besides, the early time dynamics for both thermal and MBL phases are very similar to the non-interacting dynamics we reported before. This fact is particularly evident in the thermal region, where the SIC of a large subsystem initially drops and then returns to values close to $2$. The early-time resemblance to the non-interacting case is because the density-density interaction plays no role at the beginning stage due to the Néel configuration initial states. Therefore, despite the steady-state chaotic behavior for thermal phases being predicted and shown to be identical to the random quantum circuit case in Fig. 4, the early-time information dynamics show very different characteristics.
@@
IV.3.2 Integrable interacting systems
The nature of information scrambling in generic interacting integrable systems remains mostly elusive Iyoda and Sagawa (2018); Schnaack et al. (2019); Caceffo and Alba (2023). Some of these works provide hints of information scrambling within integrable systems. We carry out the information dynamics simulation on XXZ+Z model, i.e. interacting AA model but with a uniform magnetic field (Z terms in spin language). This model is known as an interacting yet integrable system. The information dynamics and the steady values of SIC are shown in Fig. 24. Notably, the late-time behavior of the integrable system exhibits striking similarities to that of the chaotic systems, as observed in the thermal phase or random quantum circuits. Moreover, the time scale required for the integrable system to reach the steady state, which is $O(10)$, is significantly shorter than that of the thermal case, which is $O(100)$. One possible explanation for the different equilibrium time scales is that the small accessible Hilbert space in integrable systems results in faster convergence. The dimension of the Hilbert space for integrable systems is very limited due to an extensive number of effective conservation laws while thermal systems have exponentially large charge sectors for $L/2$ occupation.
@@ IV.3.2 Integrable interacting systems
IV.4 Results for topological systems
@@
The investigation on information scrambling in topological systems is limited and often restricted to the calculation of OTOC Dağ et al. (2020); Sedlmayr et al. (2023); Bin et al. (2023); Sur and Sen (2024). In contrast, we investigate the interplay between information dynamics and topology characteristics through the proxy of SIC developed in this work, which provides fresh insights into information perspectives of topological systems. To explore the effect of topological edge states, we utilize models with open boundary conditions in this subsection.
@@
Information dynamics results in topological phases $\Delta=-0.4$ and trivial phases $\Delta=0.4$ are shown in Fig. 25 and 28, respectively. Interestingly, in the topological phase, a substantial portion of the information remains trapped near the edge forever, at least for the non-interacting case here. In interacting topological systems, such information trapping behavior might be accompanied by a finite lifetime Dağ et al. (2020).
@@
On the contrary, in topological trivial phases, there is no information trapping on the edge, and the steady-state SIC curve is the same as trivial non-interacting extended states (pattern (a) in Fig. 2). It is also interesting to note that the information now propagates and reflects at both edges due to open boundary conditions. The velocity for the information spreading is consistent with the maximal speed given by the SSH Hamiltonian spectrum $\text{max}_{k}\frac{d\varepsilon_{k}}{dk}=2(1-\absolutevalue{\Delta})$ ($v_{max}=1.2$ for $\Delta=\pm 0.4$). For topological trivial systems without edge modes, the quasiparticle picture can still successfully apply to the information dynamics by considering the reflection of quasiparticles at both open boundaries Khetrapal and Pedersen (2024), the numerical results comparing the dynamics with quasiparticle picture predictions in topological trivial systems are given in Appendix F.

### v3

@@ 3.2.6 Random Floquet circuit with many-to-many encoding
In one-dimensional random Clifford Floquet circuits of Anderson localization phases, the mutual information is observed to be localized around the $L/2$ segment $E$ as shown in Fig. 16. Again, consistent with Anderson localization results with one-to-one encoding, in the thermodynamic limit, the information will be fully localized in the $L_{A}=L/2$ subsystem, and the steady state SIC gives $2$ for any $L_{A}/L>1/2$. Namely, the minimal subsystem size containing full information $MI=2$ is nearly fixed as $L_{A}=L/2+\xi$ and this critical size approaches $L_{A}/L=1/2$ in the thermodynamic limit.
@@
Figure 16: Steady-state SIC in 1D random Clifford Floquet circuit with different subsystem sizes $L_{A}$ under finite rate encoding scheme that encodes $L/2$ Bell pairs with $L/2$ ancilla qubits at the beginning. The results are averaged over 400 configurations.
4 Results for Hamiltonian dynamics
In this section, we shift our focus to Hamiltonian quench dynamics described by one-dimensional fermionic systems. We investigate information dynamics in non-equilibrium phases utilizing AA model and SSH model, respectively.
One representative example we investigated is the Aubry-André model without or with many-body interactions [84, 85, 86, 87, 88, 89]. This model is well-known for its Anderson localization (many-body localization) phase and extended (thermal) phase in non-interacting (interacting) cases. The Hamiltonian of the AA model reads as:
$H=$ $J\sum_{j=1}^{L}(c_{j+1}^{\dagger}c_{j}+h.c.)+U\sum_{j=1}^{L}n_{j}n_{j+1}$
$+$ $\sum_{j=1}^{L}V_{j}c_{j}^{\dagger}c_{j},$ (26)
We also investigate the interplay between information and topological properties [90], where SSH model is utilized [91]:
$H=-\sum_{j=1}^{L}(1-(-1)^{j}\Delta)c_{j}^{\dagger}c_{j+1}+h.c.$ (27)
@@ 4 Results for Hamiltonian dynamics
Unlike the random circuit case, where the applied local unitaries are generic without any symmetry or conservation laws, the Hamiltonians investigated in this section are U(1) symmetric and thus follow charge conservation. U(1)-symmetric evolution often gives rise to a spectrum of novel phenomena that are distinct from those in generic evolution without any conservation laws. Examples include OTOC [53, 52], Rényi entanglement growth [92, 93, 94, 95], symmetry restoration [96], and measurement-induced charge sharpening transitions [97].
@@
In this section, we focus on the one-to-one encoding scheme and von Neumann entropy based mutual information. Notably, fermionic GHZ states generated by one-to-all encoding are beyond Gaussian states and thus cannot be simulated by Gaussian state simulators for large-size systems even the evolution Hamiltonian itself is quadratic (see Appendix E). It is worth noting that in U(1) symmetric systems, Rényi entropy can show different dynamical behaviors (diffusive growth) than the von Neumann entropy (ballistic growth). We leave the information dynamics described by Rényi mutual information and Holevo information as an interesting future direction to explore.
4.1 Quasiparticle picture for SIC dynamics in non-interacting Hamiltonians
@@
The entanglement growth in noninteracting and integrable systems following a quench can be successfully explained by the intuitive quasiparticle picture [14, 17, 98, 99]. The quasiparticles of the quenched Hamiltonian are formed in pairs or multiplets and uniformly distributed in space. These quasiparticles move ballistically with the speed determined by the dispersion relations and there is no interaction between different momentum species of quasiparticles. Specifically, only quasiparticle pairs of momentum ±$\pm k$ with one residing on intervals $A$ and the other residing on its complementary $\bar{A}$ contribute to the entanglement entropy $S_{A}$ one unit $s_{k}$. Here $s_{k}$ is the entropy unit $s_{k}=-n_{k}\log n_{k}-(1-n_{k})\log(1-n_{k})$ where $n_{k}$ represents the quasiparticle density for momentum $k$. For Néel initial states and the XX Hamiltonian which is equivalent to free fermion hopping model, $n_{k}=1/2$ for all quasiparticle momentum $k$ with the velocity for quasiparticle as $v_{k}=\partial\varepsilon_{k}/\partial k=-2\sin k$ and entropy unit $s_{k}=1$ in the $\log_{2}$ base. The introduction of onsite disorder renders fermion operators of different momentum coupled, which is evident in Fourier transformed Hamiltonian in momentum space as $\sum_{k,k^{\prime}}\mu(k-k^{\prime})c^{\dagger}_{k}c_{k^{\prime}}$. Therefore, the onsite disorder leads to strong scattering between quasiparticles of different momentum, slowing down and invalidating the ballistic transport of quasiparticles.
@@
4.2 Results for non-interacting Hamiltonians
In this subsection, we focus on the results of the non-interacting AA model ($U=0$). The model in Eq. (26) has an analytical exact critical point separating the Anderson localization and extended phases at $w_{c}=J$ [84, 85].
In the absence of quasiperiodic potential, i.e. $w=0$, the results can be interpreted by the quasiparticle picture as described above. As shown in Fig. 17(b), the theoretical prediction based on the quasiparticle picture is qualitatively in agreement with the numerical results. The deviating time from the $MI=2$ plateau happens at $t^{*}=L_{A}/(2v_{max})$ where $v_{max}$ is the maximal speed of quasiparticles as $\text{max}_{k}\;2\sin k=2$. After $t^{*}$, the fastest quasiparticle from $E$ begins to leave the interval $A$, leading to the decline in SIC. Additionally, the finite size system with periodic boundary conditions renders the oscillation of SIC dynamics in the clean limit $w=0$ when the quasiparticle picture is valid without scattering between different species. The oscillation behavior is similar as reported in [100] where the quasiparticles can go around the system and get back to the interval $A$ in late times. It is important to note that the presence and absence of oscillation behaviors and the exact periodicity for the oscillation both depend on the quasiparticle dispersion relations $\epsilon_{k}$. Our results here are determined by the free fermion dispersion $\epsilon_{k}\propto\cos k$. Investigating the (non)oscillation behaviors for other dispersion relations presents a promising avenue for future research.
@@ 4.2 Results for non-interacting Hamiltonians
The late-time steady-state behavior for the SIC $I(A:R)$ is given in Fig. 19. For the clean limit $w=0$, the late-time SIC forms a straight diagonal line, indicating a uniform spatial distribution of the initial information at late times. A slightly larger $w$ will make the line bend towards the upper left. After reaching Anderson localization phases $w>w_{c}$, the SIC maintains the value $2$ for any $L_{A}/L>0$ in the thermodynamic limit. The curves for different system sizes converge in Anderson localization phases as shown in the inset of Fig. 19, reflecting the system size-independent nature of the localization length. This is similar to the Anderson localization phase in the random Clifford Floquet circuit. We further investigate the initial state dependence of SIC by simulation on dynamics from different initial states such as bipartite state and random product states as shown in Fig. 20. These results compare the late-time averaged SIC for different initial product states under the same Hamiltonian dynamics. Our findings indicate that while the numerical values of SIC can show some variation, the qualitative features, such as the spatial patterns of information distribution and the distinction between different dynamical phases, remain robustly consistent across different initial states. This supports the utility of SIC as a probe of the system’s fundamental dynamical characteristics.
@@
Figure 20: Steady-state SIC for the model defined in Eq. (26) with $J=1,U=0,L=200$ and periodic boundary conditions for different $w$. (a) Initial states for the dynamics is bipartite product state |000111..|000...111.$\rangle$| 000 … 111 . . ⟩, (b) SIC is averaged over different random product state at half-filling sector.
@@
4.3 Results for interacting Hamiltonians
@@
4.3.1 Thermal and many-body localization systems
Many-body localization is a novel non-equilibrium phase that challenges the eigenstate thermalization hypothesis [2]. Building upon Anderson localization [101], the system can evade thermalization with many-body interactions [102]. A variety of mechanisms are proposed to stabilize MBL phases, including strong random disorder [103, 104, 105, 106, 107], quasiperiodic potentials [86, 87, 88, 89, 108] and linear potentials [109, 110, 111, 112, 113]. As a fundamental basis for other emergent dynamical phases such as discrete time crystal [114, 115] and Hilbert space fragmentation [111, 112, 116], MBL has been extensively explored on various experimental platforms of programmable quantum simulators and quantum computers [117, 118, 119, 120]. In this work, we explore MBL phases through the lens of information protection and information dynamics.
Numerous studies have explored information scrambling in MBL systems, mainly using the probe of OTOC [121, 122, 123, 124, 125] and TMI [126, 127, 128]. However, these studies have not quantitatively determined the late-time redistribution of information from local perturbations. In this work, we use the newly introduced subsystem information capacity as the probe to investigate the information dynamics in MBL systems. The results are anticipated to be very different from Anderson localization cases since entanglement growth in the two phases is very different (no growth vs. logarithmic increase) [129, 130].
@@ 4.3.1 Thermal and many-body localization systems
The phase diagram of the interacting AA model is determined in Appendix F with $w_{c}\approx 1.4\pm 0.2$ for MBL transitions in our parameter regimes. The dynamical and steady-state SIC for different subsystem sizes $L_{A}$ and different potential strengths $w$ are depicted in Fig. 21 and 22, respectively. Regarding the dynamics, we observe a logarithmic decay in SIC in MBL phases, rendering a very slow saturation for the subsystem SIC as shown in Fig. 21(b). This logarithmic dynamical scaling is consistent to the logarithmic lightcone in MBL phases. Besides, the early time dynamics for both thermal and MBL phases are very similar to the non-interacting dynamics we reported before. This fact is particularly evident in the thermal region, where the SIC of a large subsystem initially drops and then returns to values close to $2$ given by information scrambling. The early-time resemblance to the non-interacting case is because the density-density interaction plays no role at the beginning stage due to the Néel configuration initial states. Therefore, despite the steady-state chaotic behavior for thermal phases being predicted and shown to be identical to the random quantum circuit case in Fig. 4, the early-time information dynamics show very different characteristics. In other words, in the thermal phase, the information within a given subsystem $L_{A}>L$ will first drop and then backflows to the region with late-time convergence toward full information access $MI=2$. Such information backflow doesn’t occur in chaotic random circuits with one-to-one encoding and is the unique early-time feature for chaotic Hamiltonian evolution.
@@
4.3.2 Integrable interacting systems
The nature of information scrambling in generic interacting integrable systems remains mostly elusive [131, 132, 133]. Some of these works provide hints of information scrambling within integrable systems. We carry out the information dynamics simulation on XXZ+Z model, i.e. interacting AA model but with a uniform magnetic field (Z terms in spin language). This model is known as an interacting yet integrable system. The information dynamics and the steady values of SIC are shown in Fig. 23. Notably, the late-time behavior of the integrable system exhibits striking similarities to that of the chaotic systems, as observed in the thermal phase or random quantum circuits. Moreover, the time scale required for the integrable system to reach the steady state, which is $O(10)$, is significantly shorter than that of the thermal case, which is $O(100)$. One possible explanation for the different equilibrium time scales is that the small accessible Hilbert space in integrable systems results in faster convergence. The dimension of the Hilbert space for integrable systems is very limited due to an extensive number of effective conservation laws while thermal systems have exponentially large charge sectors for $L/2$ occupation.
@@ 4.3.2 Integrable interacting systems
4.4 Results for topological systems
@@
The investigation on information scrambling in topological systems is limited and often restricted to the calculation of OTOC [134, 135, 136, 137]. In contrast, we investigate the interplay between information dynamics and topology characteristics through the proxy of SIC developed in this work, which provides fresh insights into information perspectives of topological systems. To explore the effect of topological edge states, we utilize models with open boundary conditions in this subsection.
@@
Information dynamics results in topological phases $\Delta=-0.4$ and trivial phases $\Delta=0.4$ are shown in Fig. 24 and 27, respectively. Interestingly, in the topological phase, a substantial portion of the information remains trapped near the edge forever, at least for the non-interacting case here. In interacting topological systems, such information trapping behavior might be accompanied by a finite lifetime [134].
@@
On the contrary, in topological trivial phases, there is no information trapping on the edge, and the steady-state SIC curve is the same as trivial non-interacting extended states (pattern (a) in Fig. 2). It is also interesting to note that the information now propagates and reflects at both edges due to open boundary conditions. The velocity for the information spreading is consistent with the maximal speed given by the SSH Hamiltonian spectrum $\text{max}_{k}\frac{d\varepsilon_{k}}{dk}=2(1-\absolutevalue{\Delta})$ ($v_{max}=1.2$ for $\Delta=\pm 0.4$). For topological trivial systems without edge modes, the quasiparticle picture can still successfully apply to the information dynamics by considering the reflection of quasiparticles at both open boundaries [138], the numerical results comparing the dynamics with quasiparticle picture predictions in topological trivial systems are given in Appendix G.